# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

### v1

@@ 3.1.1.2   Generalized Detection for GAN/VAE
Universal Detection: A primary focus has been developing methods that utilize the inherent capabilities of VLMs, such as CLIP [173], to evaluate their effectiveness across various scenarios. To enhance video content analysis, Han et al.[91] propose using the CLIP image encoder, paired with a newly designed side network-based decoder, to effectively extract both spatial and temporal cues from video clips. Unlike previous methodologies that directly utilized a frozen CLIP model, they update their CLIP-inspired model parameters by incorporating designed textual labels into the image augmentation process.
@@
Building on the multimodal capabilities, Sun et al. develop Visual-Linguistic Face Forgery Detection (VLFFD)[93], which employs CLIP to improve face forgery detection. This approach enriches training data with detailed textual prompts and fine-grained images, employing a Coarse-and-Fine Co-training (C2F) framework that leverages both visual and linguistic knowledge. VLFFD improves forgery identification and provides in-depth interpretative insights into the locations and characteristics of the forgeries, markedly expressing its explanatory power. Zhang et al. [94] reformulate the deepfake problem as a Deepfake Detection Visual Question Answering (DD-VQA) task, which involves identifying image authenticity and providing explanations. They also create a new dataset and develop a Vision and Language Transformer-based framework that effectively detects deepfakes and generates high-quality explanations.
@@
As Multimodal Large Language Models (MLLMs) have shown significant problem-solving capabilities in various vision fields, recent studies have begun exploring their application in deepfake detection. A recent study by [95] introduces the SHIELD benchmark to evaluate the effectiveness of MLLMs in face forgery detection across both visual and acoustic modalities. Tested under various conditions, including zero-shot and few-shot scenarios, MLLMs demonstrate superior capabilities in interpretability and multimodal reasoning in deepfake dtection. In parallel, Shan et al. [96] have validated the efficacy of MLLMs through extensive qualitative and quantitative experiments, highlighting the importance of effective prompt engineering to distinguish between real and AI-generated images effectively. In a recent work[97] Li et al. conduct extensive experiments to evaluate the performance of $13$ well-known language and multimodal models on the task of detecting fake images and explore how explicit reasoning impacts the accuracy of fake detection with these models.
@@
Learning Level: Advanced Loss Functions While most deepfake detection methods traditionally rely on Softmax loss, Yin et al. [98] challenge the effectiveness of softmax loss for distinguishing subtle distinctions between real and fake images. They propose a method that utilizes both angular and Euclidean spaces to enhance class separability. By decoupling the magnitude and direction of embedding vectors, they introduce margin penalties in both domains to increase the separation between classes, thereby enhancing inter-class separability and intra-class compactness. Zhang et al. [99] further the development of discriminative features using a fractional center loss, designed to cluster real faces closer within the embedding space while distinctly separating them from fake ones. In their subsequent research, Zhang et al. [100] introduce a dual-level center loss approach that incorporates both patch-level and image-level center losses. This innovative method effectively reduces intra-class distances while expanding inter-class distances, significantly enhancing the discriminative capability of feature embeddings at both the micro (patch) and macro (image) levels.
@@
Learning Level: Disentangled Representation Under the observation that representations learned by the detector contain both forgery and content information and thus remain recognizable to content clues like face identity and background, making the detector prone to overfit training data[101][102], disentanglement learning aims to decouple the learned features into task-relevant and task-irrelevant groups, while only the task-relevant group will be used for subsequent detection.
Liang et al.[101] initially explore a framework to disentangle the forgery-relevant and content-relevant features and remove content information by only employing the decoupled forgery feature for detection. The framework is devised with two encoders for content and forgery feature extraction, respectively. To ensure the effectiveness of disentanglement, a multi-task strategy and a contrastive constraint are proposed. The work of UCF[102] extends the framework by further decoupling the forgery-relevant features into manipulation-specific and manipulation-agnostic ones, targeting a more refined approach to feature disentanglement in deepfake analysis and achieving a better generalization ability across various forgery types. Lin et al.[103] also adopt this disentanglement framework with an additional focus on demographic-relevant features to improve the fairness of deepfake detection in addition to generalization ability.
@@
Unlike the aforementioned works that decompose features from a high-level semantic perspective with the aid of multi-task strategy, Ba et al.[104] leverage a disentanglement module to extract multiple non-overlapping local features with the guidance of carefully derived Local Information Loss. These local features are subsequently fused and filtered out task-irrelevant information under the supervision of Global Information Loss.
@@
Learning level: Reconstruction Researchers have innovated methods for identifying deepfakes by harnessing insights gained during the reconstruction of facial images. Khalid et al.[105] employ a One-class Variational Autoencoder (OC-VAE), trained solely on authentic images, to identify deepfakes by analyzing deviations in reconstruction scores. He et al.[106] utilize image re-synthesis techniques such as super-resolution and denoising to extract residual detection cues. Cao et al.[107] combine reconstruction with metric constraints and classification to distinguish real from forged images effectively. In MARLIN [108], Cai et al. use self-supervised learning to reconstruct accurate facial expressions from extensive online video data. These approaches leverage facial imagery reconstruction as a pivotal element for detection.
@@
Li et al.[110] develop Face X-ray, a technique based on the assumption that face manipulations involve a blending process. Their method helps differentiate real from fake images by identifying the presence and location of blending boundaries, thereby assessing image authenticity. Dang et al.[111] emphasize the importance of detecting manipulated face images and precisely localizing altered areas. They employ an attention mechanism to refine feature map processing, which significantly enhances the ability to distinguish between genuine and counterfeit faces. Shuai et al.[112] introduce a model that effectively localizes forgery areas by utilizing RGB images and Spatial Rich Mode (SRM) noise residuals. Their method integrates multi-modal and multi-scale patch features to pinpoint potential forgery areas accurately. Dong et al.[116] argue that implicit identity leakage deteriorates detection generalization, and correspondingly propose an effective ID-unaware detection model, which is guided by authenticity and manipulation localization information. In [113], Zhang et al. create EditGuard, a system designed to detect manipulated regions within digital content. EditGuard embeds imperceptible watermarks that allow for the precise decoding of tampered areas. Recent works from Hu et al.[114] and Tantaru[115] have also contributed to this direction, with Tantaru developing a weakly-supervised method that enhances the localization of manipulated areas in images produced by deep manipulation techniques.
One significant challenge in manipulation localization is the high cost associated with annotating manipulated regions, which can be both time-consuming and expertise-intensive.
@@
Cozzolino et al. [117] develop a system incorporating a facial feature extractor, a temporal network to identify biometric discrepancies, and a generative adversarial network that predicts subject-specific motions, enhancing the detection of inconsistencies in facial features. Nirkin et al. [118] analyze discrepancies between identity embeddings from the inner(target person) and outer (source person)facial regions. This comparison helps detect forgeries by highlighting inconsistencies in identity representations. Dong et al. [119] introduce the Identity Consistency Transformer, which uses identity vectors to discern differences between the inner and outer facial areas. Their model employs a consistency loss mechanism to align similar identities and differentiate dissimilar ones. Huang et al. [120] leverage implicit identity features of fake faces with two innovative loss functions: the Explicit Identity Contrast (EIC) loss, which attracts real samples to their true identities and repels fake ones, and the Implicit Identity Exploration (IIE) loss, a margin-based approach that ensures effective differentiation of genuine and manipulated faces by maintaining appropriate intra-class and inter-class distances.
This line of methods is believed to offer broad applicability across various types of manipulation, as it does not depend on the specific methods of manipulation.
@@
Woo’s Attention-based Deepfake Detection Distiller (ADD) [121] leverages high and low-quality forgery data to detect compressed deepfakes effectively. The innovative frequency attention distiller helps the student model recover lost high-frequency components, while the correlation distillation method improves pixel correlation preservation using the Sliced Wasserstein Distance metric [122].
Addressing scenarios without access to source domain data, Kim et al. introduce the Feature Representation Transfer Adaptation Learning (FReTAL) [123]. This method uses knowledge distillation to adapt detectors to new deepfake types efficiently, focusing on latent representations and prediction logits, assuming label availability for new forgeries. Similarly, the CoReD framework addresses sequential domain adaptation challenges by minimizing catastrophic forgetting within a teacher-student model setting [124]. This process ensures the knowledge preservation in both latent representation and prediction levels.
Furthermore, to improve deepfake detection across domains without needing label access for target samples, Lv et al.’s study introduces a bidirectional adaptation strategy[125]. Forward adaptation transfers detection capabilities to new forgeries via supervised training and adversarial feature adaptation, while backward adaptation uses self-distillation to further refine and maintain feature recognition and knowledge.
@@
Guo et al.[126] approach the classification of forged images by representing the forgery attributes with multiple hierarchical labels. Their method classifies forgery from a broad category (fully-synthesized vs. partial-manipulated) down to specific forgery techniques. This hierarchical interpretation aids in better understanding the nuances of the forged image and enhances the separation of real and fake regions. By predicting the fine-grained and hierarchical path, their approach provides a more detailed analysis of the attributes that define forged images.
Guarnera et al. [127] take a similar hierarchical and multi-level approach to detect and recognize deepfakes. Their method begins by classifying images as real or AI-generated, then identifies the source of synthetic images—whether from GANs or DMs—and further pinpoints the specific architecture used.
Additionally, the controllable guide-space (GS) method introduced by Guo in [128] enhances feature discrimination across forgery domains. This method employs a well-designed guide-space that effectively segregates forgery domains and accentuates the distinction between real and forged domains in a finely controlled manner, improving the generalization of forgery features.
@@ 3.1.1.3   Detection for Diffusion Model
As diffusion models continue to evolve, their potential misuse for creating counterfeit content has raised significant concerns. The distinctness of their generation process necessitates specialized detectors to discern diffusion-generated images. Given that detection for diffusion models is in its nascent stages, most works are working on images from general domain, of which a subset contains facial forgery images. Nevertheless, these proposed frameworks can be applied to face-centric synthetic images without loss of generality.
@@
The studies in [129] and [130] analyze outputs from diffusion models and concluded that, despite significant improvements in the quality of generated results, there are still identifiable clues that can distinguish these from real samples. Ricker et al. [129] extensively test whether current detection methods designed for GANs generated samples can identify images created by diffusion models. They find that while these methods initially struggled to tell DM-generated images apart from real ones, retraining them on DM images greatly improved their accuracy, even allowing them to detect images made by GANs. Corvi et al. [130] note that models like GLIDE and Stable Diffusion produce distinct spectral peaks, which can be effectively utilized to identify synthetic images. This spectral approach adds another dimension to the detection methodologies, harnessing unique signatures inherent in diffusion models.
@@
Lorenz et al. [131] introduce the use of the lightweight multi Local Intrinsic Dimensionality (multiLID) technique, initially developed for detecting adversarial examples, to automatically detect synthetic images and identify the generator networks responsible for their creation. Following a different approach, Cheng et al. [34] present the Edge Graph Regularization (EGR) method, crafted to enhance the discriminative capabilities of detectors. This technique integrates edge graphs, derived via Sobel operators, as a regularization term into the empirical risk formulation. By doing so, it encourages the model to concurrently analyze the features of both the original images and the edge graphs, thereby refining its overall analytical precision.
@@
Transitioning from intrinsic analysis to reconstruction accuracy, Wang et al. [132] introduce Diffusion Reconstruction Error (DIRE) that measures the discrepancy between an input image and its reconstructed version using a pre-trained diffusion model. Their method reveals that while images generated by diffusion models can be accurately reconstructed by the same type of models, real images do not replicate with the same fidelity. Expanding on the concept of reconstruction, Ma et al.[133] introduce the SeDID method, addressing the limitations of DIRE[132] by utilizing the intermediate steps in diffusion models. Unlike DIRE only focusing on the initial time-step information, SeDID captures critical information during the diffusion and reverse diffusion processes, potentially enhancing the accuracy of synthetic image detection. Building on this, Luo et al. [134] capitalize on conducting reconstructions within the latent space, enhancing efficiency in two key ways: it avoids the need for complete image inversion and reconstruction, and it reduces the risk of error accumulation associated with multiple inversion and reconstruction steps.

### v3

@@ 3.1.1.2   Generalized Detection for GAN/VAE
Universal Detection: A primary focus has been developing methods that utilize the inherent capabilities of VLMs, such as CLIP [192], to evaluate their effectiveness across various scenarios. To enhance video content analysis, Han et al.[102] propose using the CLIP image encoder, paired with a newly designed side network-based decoder, to effectively extract spatial and temporal cues from video clips. Unlike previous methodologies that directly utilized a frozen CLIP model, they update their CLIP-inspired model parameters by incorporating designed textual labels into the image augmentation process. In [103], Lai et al. propose the Generalized Multi-Scenario Deepfake Detection framework (GM-DF) to enhance the generalization of models trained on multiple datasets. GM-DF uses CLIP to align visual and textual features, hybrid expert modeling for domain-specific extraction, a masked image reconstruction mechanism for detailed forgeries, and a domain-aware meta-learning strategy to improve generalization.
@@
Building on the multimodal capabilities, Sun et al. develop Visual-Linguistic Face Forgery Detection (VLFFD)[105], which employs CLIP to improve face forgery detection. This approach enriches training data with detailed textual prompts and fine-grained images, employing a Coarse-and-Fine Co-training framework that leverages both visual and linguistic knowledge. VLFFD improves forgery identification and provides in-depth interpretative insights into the locations and characteristics of the forgeries, markedly expressing its explanatory power. Zhang et al. [106] reformulate the deepfake problem as a Deepfake Detection Visual Question Answering task, which involves identifying image authenticity and providing explanations. They also create a new dataset and develop a Vision and Language Transformer-based framework that effectively detects deepfakes and generates high-quality explanations.
@@
As Multimodal Large Language Models (MLLMs) have shown significant problem-solving capabilities in various vision fields, recent studies have begun to explore their application in deepfake detection. A recent study by [107] introduces the SHIELD benchmark to evaluate the effectiveness of MLLMs in face forgery detection across both visual and acoustic modalities. Tested under various conditions, including zero-shot and few-shot scenarios, MLLMs demonstrate superior capabilities in interpretability and multimodal reasoning in deepfake detection. In parallel, Shan et al. [108] have validated the efficacy of MLLMs through extensive qualitative and quantitative experiments, highlighting the importance of effective prompt engineering to distinguish between real and AI-generated images effectively. In a recent work[109], Li et al. conduct extensive experiments to evaluate the performance of $13$ well-known language and multi-modal models on the task of detecting fake images and explore how explicit reasoning impacts the accuracy of fake detection with these models.
@@
Learning Level: Advanced Loss Functions While most deepfake detection methods traditionally rely on Softmax loss, Yin et al. [110] challenge the effectiveness of softmax loss for distinguishing subtle distinctions between real and fake images. They propose a method that utilizes both angular and Euclidean spaces to enhance class separability. By decoupling the magnitude and direction of embedding vectors, they introduce margin penalties in both domains to increase the separation between classes, thereby enhancing inter-class separability and intra-class compactness. Zhang et al. [111] further the development of discriminative features using a fractional center loss, designed to cluster real faces closer within the embedding space while distinctly separating them from fake ones. In their subsequent research, Zhang et al. [112] introduce a dual-level center loss approach that incorporates both patch-level and image-level center losses. This innovative method effectively reduces intra-class distances while expanding inter-class distances, significantly enhancing the discriminative capability of feature embeddings at both the patch and image levels.
@@
Learning Level: Disentangled Representation Disentanglement learning addresses the issue that deepfake detectors may learn representations containing both forgery and content information. This mixed representation makes the detector prone to overfitting training data by recognizing content clues such as face identity and background [113] [114]. To mitigate this, disentanglement learning aims to decouple learned characteristics into task-relevant and task-irrelevant groups, ensuring that only task-relevant characteristics are used for subsequent detection.
Liang et al. [113] explore a framework to separate forgery-relevant from content-relevant features, using two dedicated encoders and employing a multi-task strategy and contrastive constraints to enhance the effectiveness of disentanglement. The UCF framework [114] further refines this approach by decoupling forgery-relevant features more distinctly, which aids in better generalization across various forgery types. Lin et al. [115] extend disentanglement learning to include demographic-relevant features, improving both fairness and generalization in deepfake detection. Similarly, Ye et al. [116] utilize an adaptive high-pass module and a two-stage training strategy to enhance the separation of forgery semantics from irrelevant content, aiming to reduce overfitting and improve generalizability. Yang et al. [117] treat different forgery methods as distinct domains and develop the Deep Information Decomposition (DID) framework, which uses attention modules to adaptively separate facial information into forgery-specific categories. This selective focus helps in enhancing detection across different datasets and forgery techniques.
@@
Unlike the aforementioned works that decompose features from a high-level semantic perspective with the aid of multi-task strategy, Ba et al.[118] leverage a disentanglement module to extract multiple non-overlapping local features with the guidance of carefully derived Local Information Loss. These local features are subsequently fused and filtered out task-irrelevant information under the supervision of introduced Global Information Loss.
@@
Learning level: Reconstruction Researchers have innovated methods for identifying deepfakes by harnessing insights gained during the reconstruction of input. Khalid et al.[119] employ a One-class Variational Autoencoder (OC-VAE), trained solely on authentic images, to identify deepfakes by analyzing deviations in reconstruction scores. He et al.[120] utilize image re-synthesis techniques such as super-resolution and denoising to extract residual detection cues. Cao et al.[121] combine reconstruction with metric constraints and classification to effectively distinguish real from forged images. In MARLIN [122], Cai et al. use self-supervised learning to reconstruct accurate facial expressions from extensive online video data. These approaches leverage facial imagery reconstruction as a pivotal element.
@@
Li et al.[123] developed Face X-ray, a technique that detects blending boundaries used in face manipulations to differentiate between genuine and altered images. Dang et al.[124] enhance the detection of manipulated face images by using an attention mechanism that refines the processing of feature maps, improving the accuracy of identifying forgeries. Similarly, Shuai et al.[125] employ RGB images and Spatial Rich Mode noise residuals to accurately localize forgery areas by integrating multimodal and multi-scale patch features. Further contributions include Dong et al.[126], who focus on mitigating implicit identity leakage to improve detection generalization with a model that emphasizes both authenticity verification and manipulation localization. Zhang et al.[127] introduce EditGuard, which embeds imperceptible watermarks to enable precise localization of tampered regions in digital content. Recently, Tantaru et al. [129] explore novel pathways in this field. Specifically, Tantaru et al. develop a weakly-supervised technique that improves the localization of manipulated areas in images created through advanced manipulation methods, further refining the ability to detect and understand the nuances of digital forgeries.
@@
Cozzolino et al. [130] develop a system incorporating a facial feature extractor, a temporal network to identify biometric discrepancies, and a generative adversarial network that predicts subject-specific motions, enhancing the detection of inconsistencies in facial features. Nirkin et al. [131] analyze discrepancies between identity embeddings from the inner(target person) and outer (source person)facial regions. This comparison helps detect forgeries by highlighting inconsistencies in identity representations. Dong et al. [132] introduce the Identity Consistency Transformer, which uses identity vectors to discern differences between the inner and outer facial areas. Their model employs a consistency loss mechanism to align similar identities and differentiate dissimilar ones. Huang et al. [133] leverage implicit identity features of fake faces with two innovative loss functions: the Explicit Identity Contrast loss, which attracts real samples to their true identities and repels fake ones, and the Implicit Identity Exploration loss, a margin-based approach that ensures effective differentiation of genuine and manipulated faces by maintaining appropriate intra-class and inter-class distances.
@@
Woo’s Attention-based Deepfake Detection Distiller [134] utilizes both high and low-quality forgery data to identify compressed deepfakes effectively. It employs a frequency attention distiller that aids the student model in recovering lost high-frequency components and a correlation distillation method that enhances pixel correlation preservation using the Sliced Wasserstein Distance metric.
In situations where access to the source domain data is restricted, Kim et al. introduce the Feature Representation Transfer Adaptation Learning (FReTAL) [136]. This method uses knowledge distillation to adapt detectors to new deepfake types efficiently, focusing on latent representations and prediction logits, assuming label availability for new forgeries. Similarly, the CoReD framework addresses sequential domain adaptation challenges by minimizing catastrophic forgetting within a teacher-student model setting [137]. This process ensures the knowledge preservation in both latent representation and prediction levels.
Furthermore, to improve deepfake detection across domains without needing label access for target samples, the study by Lv et al. [138] introduces a bidirectional adaptation strategy. Forward adaptation transfers detection capabilities to new forgeries via supervised training and adversarial feature adaptation, while backward adaptation uses self-distillation to further refine the extracted knowledge.
@@
In this line of methods, Guo et al. [139] and Guarnera et al. [140] both implement hierarchical approaches to deepfake detection, emphasizing multi-level classification to discern real from synthetic images and further classify the type and source of forgeries. Guo et al.’s method categorizes forgeries from broad groups down to specific techniques, enhancing understanding of the forgeries through a detailed hierarchical analysis. Similarly, Guarnera et al. start with basic real vs. generated classifications, then specify whether the images are from GANs or DMs, and finally identify the precise architecture used. Additionally, the work in [141] on controllable guide-space methods focuses on improving feature discrimination by finely segregating and differentiating forgery domains, thereby boosting generalization capabilities across different forgery scenarios.
@@ 3.1.1.3   Detection for Diffusion Model
As diffusion models continue to evolve, their potential misuse for creating counterfeit content has raised significant concerns. The distinctness of their generation process necessitates specialized detectors to discern diffusion-generated images. Given that detection for diffusion models is in its nascent stages, most works are working on images from general domains, of which a subset contains facial forgery images. Nevertheless, these proposed frameworks can be applied to facial forgery images without loss of generality.
@@
The studies in [142] and [143] analyze outputs from DMs and concluded that, despite significant improvements in the quality of the generated results, there are still identifiable clues that can distinguish these from the real samples. Ricker et al. [142] extensively test whether current detection methods designed for GANs generated samples can identify images created by DMs. They find that while these methods initially struggled to tell DM-generated images apart from real ones, retraining them on DM images greatly improved their accuracy, even allowing them to detect images made by GANs. Corvi et al. [143] note that models such as GLIDE [194] and Stable Diffusion [195] produce distinct spectral peaks, which can be effectively used to identify synthetic images. This spectral approach adds another dimension to detection methodologies, harnessing unique signatures inherent in diffusion models.
@@
Lorenz et al. [144] introduce the use of the lightweight multi Local Intrinsic Dimensionality technique, initially developed for detecting adversarial examples, to automatically detect synthetic images and identify the generator networks responsible for their creation. Following a different approach, Cheng et al. [34] present the Edge Graph Regularization method, crafted to enhance the discriminative capabilities of detectors. This technique integrates edge graphs, derived via Sobel operators [196], as a regularization term in the empirical risk formulation. In doing so, the model is encouraged to simultaneously analyze the features of both the original images and the edge graphs, thereby refining its overall analytical precision.
@@
Transitioning from intrinsic analysis to reconstruction accuracy, Wang et al. [145] introduce Diffusion Reconstruction Error (DIRE) that measures the discrepancy between an input image and its reconstructed version using a pre-trained diffusion model. Their method reveals that, while images generated by diffusion models can be accurately reconstructed by the same type of models, real images do not replicate with the same fidelity. Expanding on the concept of reconstruction, Ma et al.[146] introduce a method named Stepwise Error for Diffusion-generated Image Detection (SeDID), addressing the limitations of DIRE[145] by utilizing the intermediate steps in diffusion models. Unlike DIRE that only focuses on the initial time-step information, SeDID captures critical information during the diffusion and reverse diffusion processes, potentially enhancing the accuracy of synthetic image detection. Building on this, Luo et al. [147] capitalize on conducting reconstructions within the latent space, enhancing efficiency in two key ways: it avoids the need for complete image inversion and reconstruction, and it reduces the risk of error accumulation associated with multiple inversion and reconstruction steps.
3.1.1.4   Sequential Deepfake Detection
Based on the observation that the multi-step manipulation of facial components and attributes is becoming possible due to the emergence of facial editing applications, it is crucial to enable the detection of these compound deepfakes beyond one-step manipulation. In response, sequential deepfake detection[148] is proposed, aiming to detect a sequence of facial manipulations. This novel task is significantly more challenging than conventional deepfake detection, requiring a nuanced understanding to detect the facial components and attributes being forged as well as the corresponding manipulation order.
To support this investigation, Shao et al.[148] construct the first sequential deepfake dataset, Seq-DeepFake, which includes sequentially manipulated facial images and corresponding manipulation sequence annotations. They propose SeqFakeFormer, an image-to-sequence model with an image encoder for spatial correlations and a sequence decoder to capture manipulation traces. SeqFakeFormer++[149] further enhances this by incorporating image-sequence reasoning with two modules: Image Sequence Contrastive learning for aligning global image and sequence features, and Image Sequence Matching for fine-grained alignment using a cross-attention mechanism. Hong et al. [150] approach the problem with multi-label ranking for sequence prediction. They decompose the task into binary deepfake classification, manipulation localization, and manipulation order determination. They propose a contrastive multi-instance learning formulation for classification, treating an image as a bag and its patches as instances. For sequential detection, they use multi-label learning with ranking reasoning to reflect the manipulation order by multi-label scores.